# Majorana Tensor Decomposition: A unifying framework for decompositions of fermionic Hamiltonians to Linear Combination of Unitaries

### v1

@@ IV Discussion
To assess efficiency of various LCU decompositions we present quantum resource estimates for small molecules (Table 2) and 1D hydrogen chains (Fig. 5 and Table 3), computational details are provided in Appendix B. Most of LCU methods can be partitioned into three categories based on the small molecule results: 1) low T-gate count and medium 1-norms, Pauli and Pauli-OO; 2) high T-gate count and low 1-norms, AC, OO-AC, and DF; 3) medium T-gate count and high 1-norms, MTD-CP4 and MTD-MPS. The only method that combined low 1-norms and low T-gate counts is MTD-SVD.
@@
Regarding the scaling of hardness with the system size (Table 3), the only decomposition which presented a better scaling than Pauli-OO is CP4 LCU. We note that extremely large 1-norms were obtained for the MTD-CP4 tensor decomposition for the small molecules (Table 2). We attribute this to an implementation that was used for the CP4 decomposition, and having a more robust implementation will be necessary for larger systems, as discussed in Ref. 58. At the same time, we can see how the SVD decomposition gave competitive 1-norms for these molecules, having that both the SVD and the CP4 decompositions correspond to the same $L^{4}$ LCU. On the other hand, the SVD-based 1-norms for the hydrogen chains were larger than their CP4 counterparts, having that the SVD decomposition became prohibitively expensive to compute for the larger chains. Overall, these results establish the $L^{4}$ LCU decomposition as a very promising decomposition in terms of hardness; finding the corresponding fragments in a robust and efficient way is work in progress. We believe that this decomposition could be useful for hybrid LCU approaches by obtaining a few large MTD-$L^{4}$ fragments before decomposing the rest of the Hamiltonian with some other method. However, exploring these kind of hybrid LCUs is outside the scope of this work.
@@
The classical pre-processing cost: The Pauli products decomposition of a Hamiltonian has $\mathcal{O}(N^{4})$ operators. However, orbital localization schemes can significantly reduce this to $\mathcal{O}(N^{2})$ for large molecules [60]; the cost for optimized orbital localization schemes typically scales as $\mathcal{O}(N^{3})$ [31]. The sorted insertion algorithm, used in AC, has a sorting cost of $\mathcal{O}(M^{2})$, where $M$ is the number of elements in the list being sorted. This results in a decomposition cost of $\mathcal{O}(N^{8})$ for AC and $\mathcal{O}(N^{4})$ when grouping is done on localized orbitals. For the fermion-based methods, the cost for DF is dominated by the diagonalization of the reshaped $N^{2}\times N^{2}$ tensor $g_{ij;kl}$, which scales as $\mathcal{O}(N^{6})$. For the MTD-MPS, the cost for each step in the MPS factorization (see Fig. 3) goes as $\mathcal{O}(N^{5})$, thus being the overall scaling cost of this decomposition. The MTD-SVD decomposition quickly becomes prohibitively expensive to do as the system size grows, while the MTD-CP4 decomposition requires $\mathcal{O}(N^{3}W)$, where $W$ is the number of terms in the decomposition [58]. Among the algorithms presented in this work, orbital-localized AC has the best theoretical scaling, followed by MTD-MPS. However, in practice, the fermionic-based decompositions (DF and MTD-MPS) tend to have lower prefactors due to their efficient utilization of linear algebra routines. For example, in the case of the linear H30 chain, orbital-localized AC took approximately 5 minutes, while DF required around 0.1 seconds and MTD-MPS required around 0.3 seconds. These calculations were done using a single core of a 2.2 GHz 6-Core Intel i7 processor.
@@ A.2.4 MTD-$L^{4}$ decomposition
The implementation of the $\mathtt{SELECT}$ circuit for the MTD-$L^{4}$ decomposition is shown in Fig. 14. This implementation minimizes the T-gate count in exchange for a higher number of ancilla qubits: four registers are used simultaneously for loading the Givens rotation angles, minimizing the number of times QROM must be called for angle loading. The cost for the $L^{4}$ $\mathtt{SELECT}$ circuit is:
@@
The total complexity for the $L^{4}$-T $\mathtt{SELECT}$ circuit is:
@@
Figure 14: $\mathtt{SELECT}$ circuit for implementing $L^{4}$ LCU.
@@ A.2.5 Matrix Product States
The $\mathtt{SELECT}$ circuit for the MPS LCU can be seen in Fig. 15. The associated complexity components are:
@@
Figure 15: $\mathtt{SELECT}$ circuit for implementing MPS LCU.

### v2

@@ IV Discussion
To assess efficiency of various LCU decompositions we present quantum resource estimates for small molecules (Table 2) and 1D hydrogen chains (Fig. 5 and Table 3), computational details are provided in Appendix B. Most of LCU methods can be partitioned into three categories based on the small molecule results: 1) low T-gate count and medium 1-norms, Pauli and Pauli-OO; 2) high T-gate count and low 1-norms, AC, OO-AC, and DF; 3) medium T-gate count and high 1-norms, $L^{4}$-CP4 and $L^{4}$-MPS. The only method that combined low 1-norms and low T-gate counts is $L^{4}$-SVD.
@@
Regarding the scaling of hardness with the system size (Table 3), the only decomposition which presented a better scaling than Pauli-OO is $L^{4}$-CP4 LCU. We note that extremely large 1-norms were obtained for the $L^{4}$-CP4 tensor decomposition for the small molecules (Table 2). We attribute this to an implementation that was used for the CP4 decomposition, and having a more robust implementation will be necessary for larger systems, as discussed in Ref. 58. At the same time, we can see how the $L^{4}$-SVD decomposition gave competitive 1-norms for these molecules, having that both the $L^{4}$-SVD and the $L^{4}$-CP4 decompositions correspond to the same $L^{4}$ LCU. On the other hand, the SVD-based 1-norms for the hydrogen chains were larger than their CP4 counterparts, having that the SVD decomposition became prohibitively expensive to compute for the larger chains. Overall, these results establish the $L^{4}$ LCU decomposition as a very promising decomposition in terms of hardness; finding the corresponding fragments in a robust and efficient way is work in progress. We believe that this decomposition could be useful for hybrid LCU approaches by obtaining a few large MTD-$L^{4}$ fragments before decomposing the rest of the Hamiltonian with some other method. However, exploring these kind of hybrid LCUs is outside the scope of this work.
@@
The classical pre-processing cost: The Pauli products decomposition of a Hamiltonian has $\mathcal{O}(N^{4})$ operators. However, orbital localization schemes can significantly reduce this to $\mathcal{O}(N^{2})$ for large molecules [60]; the cost for optimized orbital localization schemes typically scales as $\mathcal{O}(N^{3})$ [31]. The sorted insertion algorithm, used in AC, has a sorting cost of $\mathcal{O}(M^{2})$, where $M$ is the number of elements in the list being sorted. This results in a decomposition cost of $\mathcal{O}(N^{8})$ for AC and $\mathcal{O}(N^{4})$ when grouping is done on localized orbitals. For the fermion-based methods, the cost for DF is dominated by the diagonalization of the reshaped $N^{2}\times N^{2}$ tensor $g_{ij;kl}$, which scales as $\mathcal{O}(N^{6})$. For the $L^{4}$-MPS, the cost for each step in the MPS factorization (see Fig. 3) goes as $\mathcal{O}(N^{5})$, thus being the overall scaling cost of this decomposition. The $L^{4}$-SVD decomposition quickly becomes prohibitively expensive to do as the system size grows, while the $L^{4}$-CP4 decomposition requires $\mathcal{O}(N^{3}W)$, where $W$ is the number of terms in the decomposition [58]. Among the algorithms presented in this work, orbital-localized AC has the best theoretical scaling, followed by $L^{4}$-MPS. However, in practice, the fermionic-based decompositions (DF and $L^{4}$-MPS) tend to have lower prefactors due to their efficient utilization of linear algebra routines. For example, in the case of the linear H30 chain, orbital-localized AC took approximately 5 minutes, while DF required around 0.1 seconds and $L^{4}$–MPS required around 0.3 seconds. These calculations were done using a single core of a 2.2 GHz 6-Core Intel i7 processor.
@@ A.2.4 MTD-$L^{4}$ decomposition
The implementation of the $\mathtt{SELECT}$ circuit for the MTD-$L^{4}$ decomposition is shown in Fig. 14. This implementation minimizes the T-gate count in exchange for a higher number of ancilla qubits: four registers are used simultaneously for loading the Givens rotation angles, minimizing the number of times QROM must be called for angle loading. The cost for the MTD-$L^{4}$ $\mathtt{SELECT}$ circuit is:
@@
The total complexity for the MTD-$L^{4}$ $\mathtt{SELECT}$ circuit is:
@@
Figure 14: $\mathtt{SELECT}$ circuit for implementing MTD-$L^{4}$ LCU.
@@ A.2.5 Matrix Product States
The $\mathtt{SELECT}$ circuit for the $L^{4}$-MPS can be seen in Fig. 15. The associated complexity components are:
@@
Figure 15: $\mathtt{SELECT}$ circuit for implementing $L^{4}$-MPS LCU.